OpenStudy (anonymous):

Find the slope m of the line that passes through the points (3, –8) and (–7, –4).

OpenStudy (anonymous):

well, to find the slope based on two point, you follow this formular m = \[\frac{ {y_2}-{y_1} }{ {x_2}-{x_1}}\] m is the slope so \[\frac{ -4 - 3 }{ -7-(-8) }\] = \[\frac{ -7 }{ 1 }\] = -7 your slope is -7